OpenStudy (anonymous):

Can anybody help me? If f(2x) = 2/(2+x) for all positive x, then 2f(x) = ???

OpenStudy (blockcolder):

\[f(2x)=\frac{2}{2+x}=\frac{2}{2+\frac{2x}{2}} \\f(x)=\frac{2}{2+\frac{x}{2}}\] Now whta is 2f(x)?

OpenStudy (anonymous):

looks to me like if \(f(2x)=\frac{2}{2+x}\) then \(f(x)=\frac{2}{2+\frac{x}{2}}=\frac{4}{4+x}\)

OpenStudy (anonymous):

Thank you! I wish i could give both of you medals...:(

OpenStudy (blockcolder):

I'll give one to satellite73 on your behalf. :D

OpenStudy (anonymous):

Thanks :)